OpenStudy (adamconner):

Medal and Fan!!! find the limiting reagent, mass of Al_2Cl_6 that can be produced, and mass of excess that is remained in this equation when 2.70g Al and 4.05g cl2 are mixed 2 Al + 3Cl_2 = Al_2Cl_6

OpenStudy (adamconner):

i found the limiting reagent which is the Cl_2 and the maximum Al2Cl6 that can be made is 5.08 g

OpenStudy (anonymous):

Whats the Mm of Al2cl6?

OpenStudy (adamconner):

266.66

OpenStudy (anonymous):

\[5.08g Al_2Cl_2(\frac{1 mole Al_2Cl_6}{266.7g Al_2Cl_2})(\frac{2 mole Al}{1 mole Al_2Cl_6})(\frac{27.0g Al}{1 mole Al})=X g Al\]

OpenStudy (anonymous):

2.70 g Al - X g Al = remaining reagents

OpenStudy (adamconner):

thanks